OpenStudy (logan13):

The lines s and t intersect at point R. What is the value of y? y = Lines t and s intersect at point R. Obtuse vertical angles measure seventy-nine plus y degrees and four y minus eight degrees.

rishavraj (rishavraj):

any pic???

OpenStudy (logan13):

yes

OpenStudy (logan13):

http://prntscr.com/a6ogk9

OpenStudy (alex6799):

i have no idea on this one sorry

OpenStudy (logan13):

@Agent47

rishavraj (rishavraj):

the same vertically opposite angles are equal

rishavraj (rishavraj):

so 4y - 8 = 79 + y

OpenStudy (logan13):

so 79 is the answer

rishavraj (rishavraj):

hmmmm rreally???? how??

OpenStudy (logan13):

i was asking so we - 79 from 87

rishavraj (rishavraj):

u wrote 'so 79 is the answer' ??? -_- and wht do u mean???? :/

OpenStudy (logan13):

i was aking my bad so we subtract 79 from 87

rishavraj (rishavraj):

nope why subtract...see u got 4y - 8 = 79 + y so 4y - y = 79 + 8 3y = 87

rishavraj (rishavraj):

@logan13 so whts y ??

OpenStudy (alex6799):

\[4y-8=79+y\] \[3y=87\] \[\frac{ 3y }{ 3 }=\frac{ 87 }{ 3 }\] \[y=\frac{ 3*29 }{ 3 }\] \[y=29\]

OpenStudy (logan13):

is i dont know im trying to ask how we get Y

rishavraj (rishavraj):

@alex6799 nicely done :)

OpenStudy (logan13):

wait 29 ?

rishavraj (rishavraj):

see to get y u need to solve the equation 4y - 8 = 79 + y isolate variable on one side of equality and constant on other

rishavraj (rishavraj):

and yeah its 29

OpenStudy (logan13):

this why i am failing math

OpenStudy (logan13):

i never get it

rishavraj (rishavraj):

@logan13 its all about practice....give more time to maths :))and all will be good :)

OpenStudy (alex6799):

its ok i get it and ill be more than happy to help

rishavraj (rishavraj):

@alex6799 u wrote u didnt know how to solve this one?

OpenStudy (alex6799):

at first i didn't but you put it into an equation and im very good at thos

rishavraj (rishavraj):

ooo good work then :))

OpenStudy (alex6799):

and now i know how to do it/ what to do